OpenStudy (anonymous):

An object is accelerating if there is a change in speed and/or ________. A. time B. position C. direction D. displacement

OpenStudy (anonymous):

@Angel_Halo23 @ashhhrodriguez @Elsa213 can yall help

OpenStudy (anonymous):

@Elsa213 can u help

Elsa213 (elsa213):

direction

OpenStudy (anonymous):

help

OpenStudy (anonymous):

Thanks @Elsa213

OpenStudy (anonymous):

i will need help again

OpenStudy (anonymous):

i will tag if i need yalls help ok?

Elsa213 (elsa213):

can i have medal?

Elsa213 (elsa213):

tank chu

OpenStudy (anonymous):

can yall help a few more times

OpenStudy (anonymous):

What is the acceleration of an object if it goes from a velocity of +25 m/s to rest in 5.0 s? A. –5 m/s2 B. +5 m/s2 C. –25 m/s2 D. +25 m/s2

OpenStudy (anonymous):

the question marks are supposed to be 1 - sign

OpenStudy (anonymous):

a minus

OpenStudy (anonymous):

@Elsa213

OpenStudy (anonymous):

nvm

Elsa213 (elsa213):

*2 months later* sorry idk c:

OpenStudy (anonymous):

bwahaha @elsa213

jabez177 (jabez177):

The answer for your 2nd question is A. :) @Elsa213, direct answers...

jabez177 (jabez177):

(final velocity - initial velocity) -------------------------------------- (final time - initial time) = ( 0 - 25 ) m/s / ( 5.0 - 0 ) s = -25/5 = -5 m/s^2